EVA Report – January 12th

Crew 238 EVA Report 12Jan2022

EVA #12

Author: Sionade Robinson

Purpose of EVA: To collect geologic, bacterial, and photographic samples.

Start time:1.30 pm
End time: 4.30 pm

Narrative: EVA crew visited Area D by Rovers Curiosity and Spirit. Successfully documenting photographically, the remarkable geographical features and collecting bacterial and radiation samples previously viewed to enhance findings and outputs

Destination: Area D

Coordinates (use UTM NAD27 CONUS): W4254400 by N516000

Participants: Turner, Pokrywka, Werner

Road(s) and routes per MDRS Map: MoonWalk Road up to Cow Dung Road, north to Braher Highway 1572, then west to Braher Highway to 1575

Mode of travel: Curiosity and Opportunity

Research Report – January 12th

Astronomy Report

Name: Agnieszka Pokrywka

Crew: 238

Date: 01/12/2022


Robotic Telescope Requested: Montana Learning Center

Objects Viewed: 104P Kowal, C 2019 L3 ATLAS

Problems Encountered: The observations of the two comets listed above were performed most likely with a partially cloudy sky therefore the pictures are not very clear.

I attach the remaining four observations (in the following order: Ceres_220108, IC1805_220108, 104PKowal_220112, C2019L3ATLAS_220112). These were composed out of 4 exposure files (Red, Green, Blue, Lum) without calibration. I aim to work on these after returning from the mission.

These are all observations I was ought to conduct during Crew 238 shift. The Astronomy Laptop will be placed downstairs in the Hab.


Not used.

Sol Summary – January 12th

Crew 238 Sol Summary Report 12Jan2022


Summary Title: Moonwalk

Author’s name: Sionade Robinson

Mission Status: Ongoing

Sol Activity Summary: Research work is continuing, including individual strategies for spirulina incubation. Our tubes are already visibly differentiated – judging tomorrow. This afternoon’s EVA activity was to the spectacular Area D, collecting photographic samples of this remarkable other-worldly landscape. Evening plans for further Health & Safety training, VR research, and discussions of challenges to human wellbeing in deep space-faring.

Look Ahead Plan: Research continuing. EVA request to follow

Anomalies in work: None.

Weather: Sunny, bright, and cold. No wind

Crew Physical Status: Nominal.

EVA: By Rovers to Area D for collection of photographic samples.

Reports to be filed: Sol, Operations, EVA Report, Journalist, and HSO report.

Support Requested: EVA Request

Additional Support: None

Operations Report – January 12th

Crew 238 Operations Report 12-01-2022


Name of person filing report: Simon Werner

Non-nominal systems: none

Notes on non-nominal systems: none

Spirit rover used: Yes
Hours: 151.0
Beginning charge: 100%
Ending charge: 93%
Currently charging: yes

Opportunity rover used: No
Hours: 82.6
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Curiosity rover used: Yes
Hours: 163.4
Beginning charge: 100%
Ending charge: 98%
Currently charging: yes

Perseverance rover used: No
Hours: 226
Beginning charge: unknown
Ending charge: 100%
Currently charging: Yes

General notes and comments: none

Summary of Hab operations:

WATER USE: 33 gallons
Water (static tank): 245
Water (loft tank): filled to 55 gallons
Water Meter: post pumping 01546953

Static to Loft Pump used – yes
Static tank pipe heater (on or off): on
Static tank heater (on or off) on

Toilet tank emptied: No

Summary of internet: nominal

Summary of suits and radios: nominal

Summary of GreenHab operations:

WATER USE: 10 gallons

Heater: On

Supplemental light: Yes

Harvest: no harvest today

Summary of Science Dome operations: None

Dual split: Off

Summary of RAM operations: RAM cleaned by HSO Turner.

Summary of any observatory issues: none

Summary of health and safety issues: none

Questions, concerns and requests to Mission Support:

Re-supply needed: paper towels for cleaning, trash bags. Oh, and a very terrible thing: we ran out of Coffee. We had just one open can when our rotation started.

Copyright © The Mars Society. All rights reserved. | Main Site